• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name L. Zhu

    Log entry time 15:55:57 on April25,2007

    Entry number 200017

    keyword=Day Shift Summary

    8:10 stopped the run 2783 for controlled access:
    Jack is going to look at the conditioner for the left arm.
    The Shield house temperature are: 83.9 for left and 75.5 for right.
    Jack turned the conditioner back on. (The breaker had latched.)
    9:50 continued Q^2=50 data taken with run 2784.
    10:00 Helium flow low in both Q1, then all the magnets are off. Heidi was paged and worked on this. The CHL recovery suction pressure was above 1.5atm due to the work on Hall C HMS. Asked controlled access to reset the magnets after the cryo group turned on another compressor. Bob fixed the missing TDC module (left arm R5) during the control access.
    1:00pm restarted data taking on he4 with run 2787.
    2:45pm MCC called to inform the problem with RF cavity, which was fixed in few minutes.
    2:38pm noticed that the left diple current dropped up ~10%. Reset the regulator and it came back.
    4:00pm pass to next shift with run 2796 going on.

    Run summary: Q^2=50 fm^-2
    ---------------------------
    2783: he4, 90uA, 537k
    2784: he4, 90uA, 1.08M
    2785-2786: cosmics with mangets down
    2787: he4, 90uA, 1.25M
    2788: he4, 90uA, 1.25M
    2789: very short run 23k due to beam trip.
    2790: he4, 90uA, 1.25M
    2791: he4, 90uA, 1.25M
    2792: he4, 90uA, left dipole low close to the end of run)
    2793: he4, 90uA, 1.25M
    2794: he4, 90uA, 1.25M
    2795: he4, 90uA, 1.25M
    2796: pass to next shift
    ----------------------------